If money is invested for a short period of time, simple interest is paid using the following formula.
I = Prt
P represents the money invested, or principal amount, r is the simple interest rate, and t represents the time, in years, the money is invested.
Given P = $1,000, t = 9 years, and I = $270, what is the simple interest rate?

Answer by jim_thompson5910(28598)

I = Prt

270 = 1000*r*9

270 = 1000*9r

270 = 9000r

270/9000 = r

0.03 = r

r = 0.03

So the simple interest rate is 3%
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------